OpenStudy (anonymous):

simplify 2x+10 over x^2+4x-5

OpenStudy (anonymous):

looks like this one will factor too

OpenStudy (anonymous):

\[\frac{2x+10}{x^2+4x-5}=\frac{2(x+5)}{(x+5)(x-1)}\] one more step and you are done

OpenStudy (anonymous):

is my answer 2/x-1

OpenStudy (anonymous):

Yes. Good job.

OpenStudy (anonymous):

*thumbs up*

OpenStudy (anonymous):

YEAH!! thank you!!^^

OpenStudy (anonymous):

(x+5) in the numerator and denominator divide to 1.